OpenStudy (anonymous):

You are given that P(A U B) = 0.7 and P(A U B') = 0.9 Then what is the value of P(A). A) 0.6 B)0.2 C)0.3 D)0.4

OpenStudy (mayankdevnani):

0.6

OpenStudy (anonymous):

How

OpenStudy (mayankdevnani):

P(A U B^c) = [1 - P(A U B)] + P(A). Thus, .9 = (1 - .7) + P(A) = .3 + P(A) ---> P(A) = .6

OpenStudy (mayankdevnani):

@Zekarias ok

OpenStudy (anonymous):

Ok understood